The financial manager of Haze company is thinking about introducing new equipment in next 4 years. If buying, the market price of the equipment is 2000 thousands, and the tax law requires the equipment depreciate in 10 years in straight-line method, after 4 years the residual value of the equipment is expected as 900 thousands. If leasing, the rental is 400 thousands per year. Will you advise the financial manager to buy or lease? (The tax rat of Haze Company is 30%) If you were the financial manager of leasing company, will you accepted the rental agreement? Please give the explanation (The tax rat of leasing company is 20%. Discount rate is 5%)
Answer #1
HAZE COMPANY:
1) NPV of buying = -2000000+PV of depreciation tax shield+PV of after tax residual value
= -2000000+200000*30%*(1.05^4-1)/(0.05*1.05^4)+990000/1.05^4 = $      -9,72,768
Calculation of after tax residual value:
Residual value $ 9,00,000
Original cost $      20,00,000
Depreciation for 4 years = 200000*4 = $        8,00,000
Book value $      12,00,000
Loss on sale = 1200000-900000 = $        3,00,000
Tax shield on loss at 30% $      90,000
After tax residual value [900000+90000] $ 9,90,000
2) NPV of leasing:
After tax lease rent = 400000*(1-30%) = $        2,80,000
NPV of leasing =-280000*(1.05^4-1)/(0.05*1.05^4) = $      -9,92,866
3) NAL (Net advantage of leasing) = -992866-(-972768) = -20098
As the NAL is negative, the FM should buy the new equipment. The rental
agreement should not be accepted.
4) LEASING COMPANY:
NPV of leasing = -2000000+200000*20%*(1.05^4-1)/(0.05*1.05^4)+960000/1.05^4+400000*(1-20%)*(1.05^4-1)/(0.05*1.05^4) = $            66,337
As the NPV of leasing is positive, the lease can be made.
Calculation of after tax residual value:
Residual value $ 9,00,000
Original cost $      20,00,000
Depreciation for 4 years = 200000*4 = $        8,00,000
Book value $      12,00,000
Loss on sale = 1200000-900000 = $        3,00,000
Tax shield on loss at 20% $      60,000
After tax residual value [900000+90000] $ 9,60,000
NOTE:
Discount rate for the lessee company is not given; taken as that of the leasing
company--5%.